# Solving Fredholm Integral Equations of the Second Kind via Wasserstein Gradient Flows

### v1

@@ 3.2 McKean–Vlasov SDE
The SDE (9) belongs to the class of McKean–Vlasov SDEs since the drift coefficient depends upon not only $\bm{X}_{t}$ but also its law, $\pi_{t}$. Under Assumption 1 and 2, if $\eta>0$ the drift coefficient of (9) is Lipschitz continuous and we can establish existence and uniqueness for the MKVSDE.
@@
Under Assumptions 1, 2(a) and 2(b), for any $\alpha,\eta>0$ there exists a unique strong solution to (9) for any initial condition $\bm{X}_{0}$ such that $\mathcal{L}(\bm{X}_{0})\in\mathcal{P}^{ac}_{1}(\mathbb{R}^{d})$.
@@
See Appendix C.1. ∎
The previous proposition is limited to the case where $\eta>0$. Indeed, if $\eta=0$ the drift is not Lipschitz continuous and the SDE (9) might be unstable, with solutions existing up to a (possibly small) explosion time. Using recent results from [30] we can establish that $\pi_{t}$ converges to the unique minimizer $\pi_{\alpha,\eta}^{\star}$ of $\mathcal{F}_{\alpha}^{\eta}$ when $t\to+\infty$. The presence of an interaction term $b^{\eta}$ which is not small w.r.t. $\alpha U$ prevents us from using more standard approaches (e.g. [11, 23, 41, 9]) and thus from obtaining quantitative convergence rates.
@@
Under Assumptions 1 and 2, for any $\alpha,\eta>0$, we have
@@
See Appendix C.2. ∎
The result above can be strengthened to show that, if $\alpha_{t}\to 0$ at the appropriate rate, then $\mathcal{F}^{\eta}(\pi_{t})$ converges towards the minimum of the unregularized functional (see [14, Theorem 4.1] and [45, Theorem 1]).
@@ 4 Particle System
A classical approach to circumvent this issue is to consider an interacting particle system $(\bm{X}_{t}^{1:N})_{t\geq 0}=\{(\bm{X}_{t}^{i,N})_{t\geq 0}\}_{i=1}^{N}$ for any $N\in\mathbb{N}^{\star}$ which approximates (9) and satisfies a classical SDE [42, 10]. We introduce the particle system $(\bm{X}_{t}^{1:N})_{t\geq 0}$ which satisfies the following SDE: for any $i\in\{1,\dots,N\}$, $\bm{X}_{0}^{i,N}\in\mathbb{R}^{d}$ and
@@
In the following proposition we establish a classical propagation of chaos result, showing that there exists a unique strong solution to (11) and that this solution approximates (9) for any finite time horizon.
@@
Under Assumptions 1, 2(a) and 2(b) for any $\alpha,\eta>0$ and $N\in\mathbb{N}^{\star}$ there exists a unique strong solution to (11) for any initial condition $\bm{X}_{0}^{1:N}$ such that $\mathcal{L}(\bm{X}_{0}^{1:N})\in\mathcal{P}_{1}^{ac}((\mathbb{R}^{d})^{N})$ and $\{\bm{X}_{0}^{i,N}\}_{i=1}^{N}$ is exchangeable. In addition, for any $T\geq 0$ there exists $c_{1}(T)\geq 0$ such that for any $N\in\mathbb{N}^{\star}$ and $\ell\in\{1,\dots,N\}$ if $\bm{X}_{0}^{\ell,N}=\bm{X}_{0}$ almost surely, then:
$\mathbb{E}\left[\sup_{t\in[0,T]}\left\|\bm{X}_{t}-\bm{X}_{t}^{\ell,N}\right\|^{2}\right]\leq\frac{c_{1}(T)}{\sqrt{N}}$
Proof.
See Appendix D.1. ∎
To obtain stronger convergence results for our particle system we use the dissipativity condition in Assumption 2(c). This dissipativity condition allows us to obtain the exponential ergodicity of the particle system using standard tools in the literature on convergence of SDEs. The bounds that we obtain, however, are not uniform w.r.t to the number of particles and, in particular, they depend on a constant $C_{N}\to\infty$ and a rate $\rho_{N}\to 1$ as $N\to+\infty$.
@@
Under Assumptions 1 and 2, for any $\alpha,\eta>0$ and $N\in\mathbb{N}^{\star}$ there exist $C_{N}\geq 0$ and $\rho_{N}\in[0,1)$ such that for any $x_{1}^{1:N},x_{2}^{1:N}\in(\mathbb{R}^{d})^{N}$ and $t\geq 0$
@@
where for any $x^{1:N}\in(\mathbb{R}^{d})^{N}$, $\pi_{t}^{N}(x^{1:N})$ is the distribution of $(\bm{X}_{t}^{1:N})_{t\geq 0}$ with initial condition $x^{1:N}$. In particular, (11) admits a unique invariant probability measure denoted $\pi^{N}\in\mathcal{P}_{p}((\mathbb{R}^{d})^{N})$ and for any $x^{1:N}\in(\mathbb{R}^{d})^{N}$ and $t\geq 0$
$\mathbf{W}_{1}(\pi_{t}^{N}(x^{1:N}),\pi^{N})\leq C_{N}\rho_{N}^{t}\left(\left\|x^{1:N}\right\|+\int_{\mathbb{R}^{d}}\left\|\tilde{x}\right\|\mathrm{d}\pi^{N}(\tilde{x})\right).$
Proof.
See Appendix D.2. ∎
Proposition 4 and 5 both show that the interacting particle system (11) is well behaved: Proposition 4 shows that for any finite time horizon $T$ (11) well approximates (9), whereas, Proposition 5 guarantees that (11) has a unique invariant measure for each finite $N$. The next result guarantees that the invariant measure $\pi^{N}$ of (11) converges to $\pi_{\alpha,\eta}^{\star}$.
@@
Under Assumptions 1 and 2, for any $\alpha,\eta>0$,
@@ 4.1 Euler–Maruyama Discretization
Having obtained a particle system (11) whose invariant distribution converges to the minimizer $\pi_{\alpha,\eta}^{\star}$ of (2), we now consider the numerical approximation of (11) using an Euler–Maruyama time discretization scheme. For any $N\in\mathbb{N}$, we consider the following time discretization given by $X_{0}^{1:N}\in(\mathbb{R}^{d})^{N}$ and for any $n\in\mathbb{N}$ and $k\in\{1,\dots,N\}$
$X_{n+1}^{k,N}=X_{n}^{k,N}+\gamma b(X_{n}^{k,N},\pi^{N}_{n})+\sqrt{2\gamma(\alpha+1)}Z_{n+1}^{k},$ (12)
where $\{Z_{n}^{k}\}_{k,n\in\mathbb{N}}$ is a family of independent standard Gaussian random variables, $\gamma>0$ is a stepsize, $b$ is given by
$b(x,\pi):=\int_{\mathbb{R}^{d}}b^{\eta}(x,z,\pi)\mathrm{d}\pi(z)-\alpha\nabla U(x),$
@@
Since under our assumptions the drift $b$ is Lipschitz continuous in both arguments (see Appendix C.1 for a proof), [10] guarantees that the Euler–Maruyama scheme (12) has strong order of convergence $1/2$: for any $N\in\mathbb{N}^{\star}$, $\ell\in\{1,\dots,N\}$, $\gamma>0$ and $T\geq 0$, there exists $c_{2}(T)\geq 0$, independent of $\gamma$, such that
$\mathbb{E}\left[\sup_{n\in\{0,\dots,n_{T}\}}\left\|\bm{X}_{n\gamma}^{\ell,N}-X_{n}^{\ell,N}\right\|\right]\leq c_{2}(T)\sqrt{\gamma}$ (13)
where $n_{T}=\lfloor T/\gamma\rfloor$.
For small values of $\gamma>0$ and large values of $n\gamma$ and $N\in\mathbb{N}$ we get that $\pi_{n}^{N}=(1/N)\sum_{k=1}^{N}\delta_{X_{n}^{k,N}}$ is an approximation of $\pi_{\alpha,\eta}^{\star}$. To obtain a smooth approximation, we can plug in this particle approximation into (1) and define $\hat{\pi}_{n}^{N}:\ \mathbb{R}^{d}\to\mathbb{R}$ such that for any $x\in\mathbb{R}^{d}$
$\hat{\pi}_{n}^{N}(x)=\varphi(x)+\frac{\lambda}{N}\sum_{k=1}^{N}\mathrm{k}(x,X_{n}^{k,N}),$ (14)
as suggested in [21] as a method for obtaining smooth approximations of the solution of a Fredholm equation from a particle approximation. Our final algorithm is summarized in Algorithm 1. For simplicity we assume initialization is carried out by simple random sampling from some initial distribution but other possibilities, including stratified sampling and using quasi Monte Carlo pointsets might further improve performance.
@@ 4.2 Implementation Guidelines
Notice that the drift of (9) involves nested expectations w.r.t. $\pi_{t}$. In particular, we can decompose the drift into
$b(x,\pi)$ $=\int b^{\eta}(x,z,\pi_{t})\mathrm{d}\pi_{t}(z)-\alpha\nabla U(x)$
@@
Naïve implementation of the evaluation of $b$ might lead to a $\mathcal{O}(N^{3})$ cost. However, precomputing $\pi^{N}_{n}\left[\mathrm{k}(X_{n}^{j,N},\cdot)\right]$ and storing its value for each $j=1,\dots,N$ allows computation at an $\mathcal{O}(N^{2})$ cost.
@@
The reference measure $\pi_{0}$ is the mechanism by which the choice of regularization is specified and can be chosen to impose particular properties (e.g. degree of smoothness, localized support) on $\pi$ or to favour solutions with such properties. While the presence of $\pi_{0}$ might seem arbitrary, (1) is an inverse problem, and thus regularization is required to obtain stable results. This manifests itself in the fact that $\mathcal{F}_{\alpha}^{\eta}$ admits a unique minimizer only when $\alpha>0$ (Proposition 1).
@@
This scheme corresponds to constructing the gradient flow for the functional (6), that, as previously discussed, does not necessarily lead to a unique minimizer. In addition, our proof of convergence of $\pi_{t}$ to the unique minimizer (Proposition 3) relies on the fact that $\alpha>0$ with som $\pi_{0}$ satisfying Assumption 2.
In our first experiment we show how the presence of $\pi_{0}$ stabilizes the minimizer $\pi_{\alpha,\eta}^{\star}$. See also [16, Appendix D.1] for a discussion on how $\pi_{0}$ might influence the results in a similar context.
@@
The value of $\alpha$ controls the amount of regularization introduced by the cross-entropy penalty, and, as shown by Proposition 1, when $\alpha>0$ the functional $\mathcal{F}_{\alpha}^{\eta}$ is coercive and thus admits a unique minimizer. The value of $\alpha$ could be fixed a priori, or, to guarantee a good trade-off between accuracy and regularization, it could be selected by cross-validation as suggested in [16, 54, 3] for Fredholm integral equations of the first kind.
@@
$\mathbb{E}\left[\sup_{n\in\{0,\dots,n_{T}\}}\left\|\bm{X}_{n\gamma}^{\star}-X_{n}^{\ell,N}\right\|\right]\leq\frac{c_{1}(T)}{\sqrt{N}}+c_{2}(T)\sqrt{\gamma}$ (16)
Choosing $N$ amounts to the classical task of selecting an appropriate sample size for Monte Carlo approximations, while the choice of $\gamma$ corresponds to the specification of a timescale on which to discretize a continuous time (stochastic) process; hence, one can exploit the vast literature on Monte Carlo methods and discrete time approximations of SDEs to select these values [37].
Given a fixed computational budget $B$, and observing that the cost of running Algorithm 1 is $\mathcal{O}(N^{2}/\gamma)$, we can minimize the r.h.s. of (16) subject to $B=N^{2}/\gamma$ and obtain, via a simple Lagrange multiplier argument, that the optimal values of $N,\gamma$ are given by $N=B^{1/3}[c_{1}(T)/(2c_{2}(T))]^{2/3}$ and $\gamma=N^{2}/B=(c_{1}(T)/2c_{2}(T))^{4/3}B^{-1/3}$. This suggests that, optimally, $N$ and $1/\gamma$ both scale with the cube root of the computational effort available—suggesting setting $\gamma=O(N^{-1})$, which results in an error decaying at the usual Monte Carlo rate of $O(N^{-1/2})$ with an overall cost scaling as $O(N^{3})$.

### v2

@@ 3.2 McKean–Vlasov SDE
Comparing (10) with the drift of the MKVSDE derived in [16, Eq. (13)] for equations of the first kind we find that the dependence of the dynamics (9) on $\pi$ is considerably more complicated. In particular, the kernel $\mathrm{k}$ needs to be integrated against $\pi$ twice (one w.r.t. the first component and one w.r.t. the second) and $b^{\eta}$ in (10) is then integrated w.r.t. to obtain the drift in (9). This is not the case for the MKVSDE derived in [16] in which $\pi$ only appears once in the definition of the drift. The presence of nested integrals w.r.t $\pi$ complicates both the analysis (e.g. establishing the Lipschitz continuity of the drift needed for the proof of Proposition 3 below requires extra care) and the numerical approximation: naïve approximations of (10) obtained by replacing $\pi$ with an empirical measure $\pi^{N}$ can lead to a cubic cost in the number of particles as discussed in Section 4.2.
The SDE (9) belongs to the class of McKean–Vlasov SDEs as the drift coefficient depends upon not only $\bm{X}_{t}$ but also its law, $\pi_{t}$. Under Assumption 1 and 3, if $\eta>0$ the drift of (9) is Lipschitz continuous and we can establish existence and uniqueness for the MKVSDE.
@@
Under Assumptions 1, 3(a) and 3(b), for any $\alpha,\eta>0$ there exists a unique strong solution to (9) for any initial condition $\bm{X}_{0}$ such that $\mathcal{L}(\bm{X}_{0})\in\mathcal{P}^{ac}_{1}(\mathbb{R}^{d})$.
@@
See Appendix D.1. ∎
The previous proposition is limited to the case where $\eta>0$. Indeed, if $\eta=0$ the drift is not Lipschitz continuous and the SDE (9) might be unstable, with solutions existing up to a (possibly small) explosion time. As highlighted before, showing that the drift of (9) is Lipschitz continuous requires extra care compared to [16, Proposition 5]: not only one needs to show that $b^{\eta}$ is Lipschitz continuous in $x$ as for [16, Proposition 5], but also that $b^{\eta}$ is Lipschitz continuous in $z$ due to the presence of the integral w.r.t. $\pi$ in (9).
Using recent results from [32] we can establish that $\pi_{t}$ converges to the unique minimizer $\pi_{\alpha,\eta}^{\star}$ of $\mathcal{F}_{\alpha}^{\eta}$ when $t\to+\infty$. The presence of an interaction term $b^{\eta}$ which is not small w.r.t. $\alpha U$ prevents us from using more standard approaches (e.g. [11, 25, 43, 9]) and thus from obtaining quantitative convergence rates.
@@
Under Assumptions 1 and 3, for any $\alpha,\eta>0$, we have
@@
See Appendix D.2. ∎
The previous result deals with the behaviour of $\pi_{t}$ for fixed $\alpha,\eta$. Under our assumptions, [14, Thm. 4.1] and [47, Thm. 1] guarantee that $\mathcal{F}^{\eta}(\pi_{t})-\inf\mathcal{F}^{\eta}\to 0$ provided that the parameter $\alpha$ decays at rate $1/\log t$.
@@ 4 Particle System
A classical approach to circumvent this issue is to consider an interacting particle system $(\bm{X}_{t}^{1:N})_{t\geq 0}=\{(\bm{X}_{t}^{i,N})_{t\geq 0}\}_{i=1}^{N}$ which approximates (9) for any $N\in\mathbb{N}^{\star}$ and satisfies a classical SDE [44, 10]. We introduce the particle system $(\bm{X}_{t}^{1:N})_{t\geq 0}$ which satisfies the following SDE, for any $i\in\{1,\dots,N\}$, $\bm{X}_{0}^{i,N}\in\mathbb{R}^{d}$:
@@
In the following proposition we establish a propagation of chaos result, showing that there exists a unique strong solution to (11) and that this solution approximates (9) for any finite time horizon. The proof is classical, see, e.g. [16, Prop. 7], and is given in the supplement for completeness (see E.1).
@@
Under Assumptions 1, 3(a) and 3(b) for any $\alpha,\eta>0$ and $N\in\mathbb{N}^{\star}$ there exists a unique strong solution to (11) for any initial condition $\bm{X}_{0}^{1:N}$ such that $\mathcal{L}(\bm{X}_{0}^{1:N})\in\mathcal{P}_{1}^{ac}((\mathbb{R}^{d})^{N})$ and $\{\bm{X}_{0}^{i,N}\}_{i=1}^{N}$ is exchangeable. In addition, for any $T\geq 0$ there exists $c_{1}(T)\geq 0$ such that for any $N\in\mathbb{N}^{\star}$ and $\ell\in\{1,\dots,N\}$ if $\bm{X}_{0}^{\ell,N}=\bm{X}_{0}$ almost surely, then:
$\mathbb{E}\left[\sup_{t\in[0,T]}\left\|\bm{X}_{t}-\bm{X}_{t}^{\ell,N}\right\|^{2}\right]^{1/2}\leq\frac{c_{1}(T)}{\sqrt{N}}$
To obtain stronger convergence results for our particle system we use the dissipativity condition in Assumption 3(c) which allows us to obtain the exponential ergodicity of the particle system using standard tools.The bounds that we obtain, however, are not uniform w.r.t to the number of particles and, in particular, they depend on a constant $C_{N}\to\infty$ and a rate $\rho_{N}\to 1$ as $N\to+\infty$. The proof (see E.2) follows the argument of [16, Prop. 8].
@@
Under Assumptions 1 and 3, for any $\alpha,\eta>0$ and $N\in\mathbb{N}^{\star}$ there exist $C_{N}\geq 0$ and $\rho_{N}\in[0,1)$ such that for any $x_{1}^{1:N},x_{2}^{1:N}\in(\mathbb{R}^{d})^{N}$ and $t\geq 0$
@@
where for any $x^{1:N}\in(\mathbb{R}^{d})^{N}$, $\pi_{t}^{N}(x^{1:N})$ is the distribution of $(\bm{X}_{t}^{1:N})_{t\geq 0}$ with initial condition $x^{1:N}$. In particular, (11) admits a unique invariant probability measure, $\pi^{N}\in\mathcal{P}_{1}((\mathbb{R}^{d})^{N})$, and for any $x^{1:N}\in(\mathbb{R}^{d})^{N}$ and $t\geq 0$
[…]
Proposition 5 and 6 both show that the interacting particle system (11) is well behaved: Proposition 5 shows that for any finite time horizon $T$ (11) well approximates (9), whereas, Proposition 6 guarantees that (11) has a unique invariant measure for each finite $N$. The next result guarantees that the invariant measure $\pi^{N}$ of (11) converges to $\pi_{\alpha,\eta}^{\star}$. The proof follows the same argument of [16, Prop. 9], and is given in E.3 for completness.
@@
Under Assumptions 1 and 3, for any $\alpha,\eta>0$,
@@ 4.1 Euler–Maruyama Discretization
Having obtained a particle system (11) whose invariant distribution converges to the minimizer $\pi_{\alpha,\eta}^{\star}$ of (2), we now consider the numerical approximation of (11) using an Euler–Maruyama time discretization scheme. For any $N\in\mathbb{N}$, we consider the following time discretization given by $X_{0}^{1:N}\in(\mathbb{R}^{d})^{N}$ and
$\forall n\in\mathbb{N},k\in\{1,\ldots,N\}:\qquad X_{n+1}^{k,N}=X_{n}^{k,N}+\gamma b(X_{n}^{k,N},\pi^{N}_{n})+\sqrt{2\gamma(\alpha+1)}Z_{n+1}^{k},$ (12)
where $\{Z_{n}^{k}\}_{k,n\in\mathbb{N}}$ is a family of independent standard Gaussian random variables, $\gamma>0$ is a stepsize, $b$ is
$b(x,\pi):=\int_{\mathbb{R}^{d}}b^{\eta}(x,z,\pi)\mathrm{d}\pi(z)-\alpha\nabla U(x),$ (13)
@@
Since under our assumptions the drift $b$ is Lipschitz continuous in both arguments (see Appendix D.1 for a proof), [10] guarantees that the Euler–Maruyama scheme (12) has strong order of convergence $1/2$: for any $N\in\mathbb{N}^{\star}$, $\ell\in\{1,\dots,N\}$, $\gamma>0$ and $T\geq 0$, there exists $c_{2}(T)\geq 0$, independent of $\gamma$, such that for $n_{T}=\lfloor T/\gamma\rfloor$:
$\mathbb{E}\left[\sup_{n\in\{0,\dots,n_{T}\}}\left\|\bm{X}_{n\gamma}^{\ell,N}-X_{n}^{\ell,N}\right\|\right]\leq c_{2}(T)\sqrt{\gamma}.$ (14)
For simplicity, we use an Euler–Maruyama time discretization instead of the tamed one proposed in [16]; the convergence rate in (14) follows from [10], to obtain a stronger order of convergence one would need additional control on the drift of (11) and its Lions derivatives as shown in [16, Appendix C].
For small values of $\gamma>0$ and large values of $n\gamma$ and $N\in\mathbb{N}$ we get that $\pi_{n}^{N}=(1/N)\sum_{k=1}^{N}\delta_{X_{n}^{k,N}}$ is an approximation of $\pi_{\alpha,\eta}^{\star}$. To obtain a smooth approximation, we can plug in this particle approximation into (1) and define $\hat{\pi}_{n}^{N}:\ \mathbb{R}^{d}\to\mathbb{R}$ via:
$\forall x\in\mathbb{R}^{d}:\qquad\hat{\pi}_{n}^{N}(x)=\varphi(x)+\frac{\lambda}{N}\sum_{k=1}^{N}\mathrm{k}(x,X_{n}^{k,N}),$ (15)
as suggested in [23] as a method for obtaining smooth approximations of the solution of a Fredholm equation of the second kind from a particle approximation.
Our final algorithm is summarized in Algorithm 1. For simplicity we assume initialization is carried out by simple random sampling from some initial distribution but other possibilities, including stratified sampling and using quasi Monte Carlo pointsets might further improve performance.
@@ 4.2 Implementation Guidelines
Notice that the drift of (9) involves nested expectations w.r.t. $\pi_{t}$ contrary to the drift of the MKVSDE obtained for equations of the first kind in [16]. In particular, we can decompose the drift into
$b(x,\pi)$ $=\int b^{\eta}(x,z,\pi)\mathrm{d}\pi(z)-\alpha\nabla U(x)$
@@
Naïve implementation of the evaluation of $b$ can lead to $\mathcal{O}(N^{3})$ cost; precomputing $\pi^{N}_{n}\left[\mathrm{k}(X_{n}^{j,N},\cdot)\right]$ and storing its value for each $j$ allows computation at $\mathcal{O}(N^{2})$ cost.
Particle systems such as this also lend themselves to natural parallel implementation. On appropriate shared-memory architectures one could parallelise both the pre-computation of $\pi^{N}_{n}\left[\mathrm{k}(X_{n}^{j,N},\cdot)\right]$ and then the propagation of the particles themselves leaving an $\mathcal{O}(N)$ cost for direct parallel implementations with width $N$; further performance improvements could be obtained by using fork-and-join techniques to carry out the computations associated with individual particles.
@@
The reference measure $\pi_{0}$ is the mechanism by which regularization is specified and can be chosen to impose particular properties (e.g. degree of smoothness, localized support) on $\pi$ or to favour solutions with such properties. While the presence of $\pi_{0}$ might seem arbitrary, (1) is an inverse problem, and thus regularization is required to obtain stable results. This manifests itself in the fact that $\mathcal{F}_{\alpha}^{\eta}$ admits a unique minimizer only when $\alpha>0$ (Proposition 1).
@@
This scheme corresponds to constructing the gradient flow for the functional (6), that, as previously discussed, does not necessarily lead to a unique minimizer. In addition, our proof of convergence of $\pi_{t}$ to the unique minimizer (Proposition 4) relies on the fact that $\alpha>0$ with some $\pi_{0}$ satisfying Assumption 3.
The first experiment shows that the presence of $\pi_{0}$ stabilizes the minimizer $\pi_{\alpha,\eta}^{\star}$. See also [16, App. D.1] for a discussion of the impact of $\pi_{0}$ in a similar context.
@@
The value of $\alpha$ controls the amount of regularization introduced by the cross-entropy penalty, and, as shown by Proposition 1, when $\alpha>0$ the functional $\mathcal{F}_{\alpha}^{\eta}$ is coercive and thus admits a unique minimizer. The value of $\alpha$ could be fixed a priori, or, to guarantee a good trade-off between accuracy and regularization, it could be selected by cross-validation as suggested in [16, 58, 3] for integral equations of the first kind.
@@
$\mathbb{E}\left[\sup_{n\in\{0,\dots,n_{T}\}}\left\|\bm{X}_{n\gamma}^{\star}-X_{n}^{\ell,N}\right\|\right]\leq\frac{c_{1}(T)}{\sqrt{N}}+c_{2}(T)\sqrt{\gamma}.$ (17)
This global error estimate differs from the one in [16, Eq. (19)]: for equations of the second kind, $\pi$ appears on both sides of (1) and thus no observed data is available (leading to the lack of the $\gamma m^{-1/2}$ term in [16, Eq. (19)]); the different analysis of the time discretization scheme discussed below (14) leads to the different rate w.r.t. $\gamma$.
Choosing $N$ amounts to the classical task of selecting an appropriate sample size for Monte Carlo approximations, while the choice of $\gamma$ corresponds to the specification of a timescale on which to discretize a (stochastic) process; hence, one can exploit the vast literature on Monte Carlo methods and discrete time approximations of SDEs to select these values [39].
Given a fixed computational budget $B$, as the cost of running Algorithm 1 is $\mathcal{O}(N^{2}/\gamma)$, we can minimize the r.h.s. of (17) subject to $B=N^{2}/\gamma$ and obtain, via a simple Lagrange multiplier argument, that the optimal values of $N,\gamma$ are $N=B^{1/3}[c_{1}(T)/(2c_{2}(T))]^{2/3}$ and $\gamma=N^{2}/B=(c_{1}(T)/2c_{2}(T))^{4/3}B^{-1/3}$. This suggests that, optimally, $N$ and $1/\gamma$ both scale with the cube root of the computational effort available—suggesting setting $\gamma=O(N^{-1})$, which results in an error decaying at the usual Monte Carlo rate of $O(N^{-1/2})$ with an overall cost scaling as $O(N^{3})$.